OpenStudy (anonymous):

sin(x+3pi/2)cosx....simplify

OpenStudy (anonymous):

sin(x+3pi/2) = -sin(x) -sin(x)cos(x) = -(1/2)sin(2x) (trig identity -- 2sinxcosx = sin(2x))